OpenStudy (dls):

HELP!! Co-ordinate Geometry~ The equation of the straight line which passes through the point(-4,3) such that the portion of the life between the axes is divided internally by the point in the ratio 5:3 is ?

OpenStudy (dls):

 is the diagram correct?

OpenStudy (dls):

@Chlorophyll @hba

OpenStudy (dls):

@hartnn

hartnn (hartnn):

so, how would u use this information 'divided internally by the point in the ratio 5:3'

OpenStudy (dls):

\[\LARGE 3=\frac{3y}{8} => y=8\]

OpenStudy (dls):

similary.. \[\LARGE -4=\frac{5x}{8}=> x=-32/5\]

OpenStudy (dls):

\[\LARGE m = \frac{8 \times 5}{-32} => \frac{-5}{4}\]

hartnn (hartnn):

are u sure ? its 3=3y/8 or 3=5y/8 ??

OpenStudy (dls):

former

hartnn (hartnn):

then u are correct. so whats the equation? u got slope, and points on the line...

OpenStudy (dls):

\[\LARGE y-8=\frac{-5}{4}(x+\frac{32}{5}\]

hartnn (hartnn):

no.! which point did u use ?

hartnn (hartnn):

u have 3 points now -4,3 0,8 -32/5,0 i suggest u use 0,8

OpenStudy (dls):

y=mx+c 3=-5+c c=8 y=-5/4x+c

hartnn (hartnn):

yeah

OpenStudy (dls):

\[\LARGE y= \frac{-5}{4x}+8\]

OpenStudy (dls):

x in numerator^^

hartnn (hartnn):

c is y-intercept, and u already found it as 8 as(8,0) and yes.

OpenStudy (dls):

\[4y+5x-8=0?\]

hartnn (hartnn):

-8 or -32 ?

OpenStudy (hba):

y-intercept x=0

OpenStudy (dls):

-32 sorry^^

hartnn (hartnn):

then ok.

OpenStudy (dls):

answer is wrong though :P

hartnn (hartnn):

whats the answer ?

OpenStudy (dls):

\[\LARGE 9x-20y+96=0 \] /S

OpenStudy (hba):

x:y=5:3 or x:y=3:5

OpenStudy (hba):

I remember doing this 2 years ago :(

OpenStudy (dls):

where did we go wrong >.<

hartnn (hartnn):

i don't even remember doing this a decade ago :P

OpenStudy (hba):

lol

hartnn (hartnn):

well, did u try 5=3y/8 ?

hartnn (hartnn):

3=5y/8

OpenStudy (dls):

y=24/5 ?

OpenStudy (dls):

x=-32/5

hartnn (hartnn):

just try it...

OpenStudy (dls):

m=-3/2

hartnn (hartnn):

is your x correct ?

OpenStudy (dls):

100%

hartnn (hartnn):

coz i just got x=-32/3, don't know how

OpenStudy (dls):

x=(m2x1+m1x2)/m1+m2

OpenStudy (dls):

-4=-5x/8

OpenStudy (dls):

+5x*

hartnn (hartnn):

then 4y+5x−8=0 should be correct

OpenStudy (dls):

but its not

OpenStudy (dls):

i got that before too

hartnn (hartnn):

i mean 4y+5x−32=0 u sseing the right answer ?

OpenStudy (dls):

yes :/ lol i have options too

OpenStudy (dls):

\[9x-20y+96=0\] \[9x-20y=24\] \[20x-9y+53=0\] \[N/T\]

OpenStudy (dls):

PERHAPS,we interpreted the question wrong?

hartnn (hartnn):

i think that interpretation and diagram is correct.

OpenStudy (dls):

well..

OpenStudy (shubhamsrg):

here is my attempt : let the intercepts be a and b such that eqn is x/a + y/b =1 hence we are given (-4,3) divides (a,0) and (0,b) in the ratio 5/3 => -4 = 3a/8 a = -32/3 =>3 = 5b/8 b=24/5 then ans is x/(-32/3) + y/(24/5) =1 on simplification, am easily getting 9x -20y + 96 =0 o.O

OpenStudy (dls):

lol o.o